OpenStudy (anonymous):

I need help transfoming this equation into slope-intercept form. 3/5x-3/2y=-6

OpenStudy (anonymous):

slope intercept form of an equation follow the form "y=mx+b" where m is the slope and b is the intercept. To transform your equation to such form you need to isolate "y". subtract (3/5)x from both sides: (-3/2)y = -6-(3/5)x multiply by (-2/3) on both sides:(-3/2)(-2/3)y = (-6)(-2/3) -(3/5)(-2/3)x simplify: y = 4 + (2/5)x